OpenStudy (anonymous):

What valur completes the square for x^2+2x?

OpenStudy (anonymous):

x^2+2x+1 (x+1)(x+1)

OpenStudy (wolf1728):

x^2+2x = 0 To complete the square get coefficient of x and divide it by 2 then square it 2 divided by 2 =1 then squared = 1 Add to both sides of the equation: x^2+2x +1 = 1 (x+1)*(x+1) = 1 x1=0 x2=-2 To read about completing the square, go here: http://www.1728.org/quadr2.htm

Directrix (directrix):

@shahmad As I understand the question, the task is to give the number required to make x^2+2x into a perfect square trinomial. As @wolf1728 stated, do this to find that number: "To complete the square get coefficient of x and divide it by 2 then square it 2 divided by 2 =1 then squared = 1" So, the answer is simply 1. --> There is no equation to solve; there was no equation given in the posted problem.